The density of chromium metal is 7.14 g/cm3. A cube of chromium has a mass of 32.00 kilograms. Calculate the length of each side of the cube in inches.

1 inch = 2.54 cm exactly.
1
Expert's answer
2019-09-16T06:07:00-0400

p=m/V V=m/p = 32000 / 7.14 = 4481.79272 cm3

V= a3 a= V^(1/3)= 4481.79272^(1/3) = 16.4873398 cm